OpenStudy (anonymous):

The original purchase price of a car is 12,000 Each year its value depreciates by 5% Three years after its purchase, what is the value of the car?

OpenStudy (tkhunny):

What is the value after one year?

OpenStudy (harsha19111999):

5% for a year is 15% for 3 years. So, 12000*15/100 = ??

OpenStudy (anonymous):

1800

OpenStudy (shamim):

u can also use the formula of simple interest I=Pnr

OpenStudy (harsha19111999):

12000-1800= 10200

OpenStudy (anonymous):

mmmmhh ok

OpenStudy (harsha19111999):

Therefore value of car= 10200

OpenStudy (anonymous):

oh ok then mult. by three ...?

OpenStudy (harsha19111999):

Sorry I didn't get u

OpenStudy (anonymous):

10200 * 3 cuzz its three years ..

OpenStudy (harsha19111999):

no

OpenStudy (harsha19111999):

you have calculated the depreciation for 3 years already

OpenStudy (harsha19111999):

10200 is the final answer

OpenStudy (anonymous):

ohh ok

OpenStudy (harsha19111999):

:)

OpenStudy (anonymous):

thanks

OpenStudy (harsha19111999):

ur welcome

OpenStudy (tkhunny):

?? Why are we suing a Simple Interest formula? Notice: Original Price: 12000 First Depreciation: .05*12000 = 600 Percent Depreciation: 600/12000 = 0.05 = 5% Value after One Year: 12000 - 600 = 11400 If we use the same DOLLAR depreciation, look what happens. Second Depreciation: 600 Percent Depreciation: 600/11400 = 0.052632 = 5.2632% This violates the premise of the problem statement. The depreciation should be 5%, not 5.2632%! If something depreciates 5% each year, then each year the value is decreased 5%. This is not a constant value. Decreasing 5% suggests that there is 95% left. 100% - 5% = 95% We should be doing this: Value(t) = 12000*0.95^t Value(0) = 12000*0.95^0 = 12000 -- As required Value(1) = 12000*0.95^1 = 11,400 -- As Before Value(2) = 12000*0.95^2 = 10,830 Value(3) = 12000*0.95^3 = 10,288.5

OpenStudy (harsha19111999):

Hey u r using compound interest which is not needed here

OpenStudy (tkhunny):

I explained the need for the process. If you are okay using 5.26% when the problem statement says 5%, go right ahead and do it that way. If, however, you're happier using 5% when the problem statement says 5%, you should read my presentation more carefully. Using the method you prescribed, what is the depreciated value after 21 years?